# The JADES Origins Field: A New JWST Deep Field in the JADES Second NIRCam Data Release

### v1

@@ Abstract
We summarize the properties and initial data release of the JADES Origins Field (JOF), which will soon be the deepest imaging field yet observed with the James Webb Space Telescope (JWST). This field falls within the GOODS-S region about 8’ south-west of the Hubble Ultra Deep Field (HUDF), where it was formed initially in Cycle 1 as a parallel field of HUDF spectroscopic observations within the JWST Advanced Deep Extragalactic Survey (JADES). This imaging will be greatly extended in Cycle 2 program 3215, which will observe the JOF for 5 days in six medium-band filters, seeking robust candidates for galaxies. This program will also include ultra-deep parallel NIRSpec spectroscopy (up to 104 hours on-source, summing over the dispersion modes) on the HUDF. Cycle 3 observations from program 4540 will add 20 hours of NIRCam slitless spectroscopy to the JOF. With these three campaigns, the JOF will be observed for 380 open-shutter hours with NIRCam using 15 imaging filters and 2 grism bandpasses. Further, parts of the JOF have deep 43 hr MIRI observations in F770W. Taken together, the JOF will soon be one of the most compelling deep fields available with JWST and a powerful window into the early Universe. This paper presents the second data release from JADES, featuring the imaging and catalogs from the year 1 JOF observations.
@@ 1 Introduction
A primary goal of the James Webb Space Telescope (JWST) is to study galaxies to the earliest observable epochs, when the first massive assemblies of gas and stars set in motion the long play of galaxy evolution. Precisely timing when the first galaxies formed after the Big Bang is a key goal of observational cosmology and one of the last great unknowns in our history of the Universe: When and how did the first galaxies form? The primordial star formation and the formation of galaxies in the first 300 Myr (redshift ) after the Big Bang is a sensitive probe of the physics governing the conversion of gas into stars at early times and different dark matter models (e.g., Behroozi & Silk, 2015; Dayal & Ferrara, 2018; Boylan-Kolchin, 2023; Kannan et al., 2023). Additionally, these first, unexplored galaxies might harbor surprises such as metal-free stellar populations (Population III) or rapidly accreting young black hole seeds, thereby shedding light onto possible pathways to the formation of the billion-solar-mass black holes observed at (e.g., Bromm & Yoshida, 2011; Xu et al., 2013; Smith & Bromm, 2019; Vikaeus et al., 2022; Volonteri et al., 2021; Trinca et al., 2023; Schneider et al., 2023). Deep fields, both in imaging and associated spectroscopy, are a key opportunity for this work, allowing the astronomy community to coordinate its multi-wavelength resources to best advantage.
@@
Cycle 1 observations with JWST succeeded quickly in pushing the redshift frontier beyond what had been seen with HST. Relying on the complete scattering of photons shortward of Ly 1216Å by the intergalactic medium (IGM), many F115W dropouts at and even some F150W dropouts at were detected as photometric candidates in numerous fields (e.g., Bouwens et al., 2022; Castellano et al., 2022; Finkelstein et al., 2022; Harikane et al., 2023a; Naidu et al., 2022b; Adams et al., 2023; Austin et al., 2023; Atek et al., 2023; Donnan et al., 2023; Franco et al., 2023; Hainline et al., 2023; Wang et al., 2023a; Whitler et al., 2023; Yan et al., 2023). As of this writing, numerous galaxies at , including two at , have been spectroscopically confirmed with long NIRSpec exposures (Curtis-Lake et al., 2023; Heintz et al., 2023; Morishita & Stiavelli, 2023; Roberts-Borsani et al., 2023; Tang et al., 2023; Arrabal Haro et al., 2023b; Bunker et al., 2023b; Hsiao et al., 2023; Sanders et al., 2023b; Wang et al., 2023b; Fujimoto et al., 2023b), along with a larger number of galaxies out to where the much brighter [OIII] and H lines are available in the NIRSpec bands.
@@
These early results have demonstrated not only JWST’s unrivalled capabilities in successfully exploring the early Universe, but also its potential for pushing the redshift frontier even further. Indeed, it is evident that JWST—with its excellent design and on-orbit performance (Rigby et al., 2022)—is fully capable of probing back to even earlier cosmic times. Within this context, we note that the Cycle 1 programs were designed to go deep in F115W and F150W, not in F200W. The few hints of any F200W dropouts are uncertain, potentially mimicked by –5 galaxies with dust-reddened continua and well-placed strong rest-optical emission lines (Fujimoto et al., 2023a; Naidu et al., 2022a; Zavala et al., 2022; Arrabal Haro et al., 2023a).
@@
We describe here the construction of a new deep field, the JADES Origins Field (JOF), that aims to push our exploration of early galaxies further than ever before possible. This field was initiated by the JWST Advanced Deep Extragalactic Survey (JADES; Eisenstein et al., 2023) as JWST NIRCam coordinated parallel observations to a NIRSpec multi-object spectroscopy (MOS) pointing focused on the Hubble Ultra Deep Field (HUDF; Beckwith et al., 2006) within the GOODS-S field (Giavalisco et al., 2004). These Cycle 1 observations of JADES were described in Eisenstein et al. (2023) and Bunker et al. (2023a). New data in Cycle 2 and 3 will bring this field to nearly 10 mission days of NIRCam imaging and slitless spectroscopy in addition to over 7 mission days of deep MIRI imaging, making it the deepest JWST imaging yet achieved. The geometry of this field is well suited for future JWST work, as the parallels can span between the JADES Deep mosaic around the HUDF and the JOF at observing windows that have low zodiacal background. As a start on this, NIRSpec MOS work in the JOF will be conducted as part of JADES in late 2023.
@@
In this paper, we detail the scientific design and implementation of the Cycles 2 and 3 observations. Relative to other deep fields with JWST, this combined data set features very deep medium-band imaging in a total of 15 NIRCam filters. The medium-band imaging is motivated by the goal of having finer sampling of the spectral energy distribution for searches for Ly dropouts at . As we argue in § 3 (and as mentioned above), broad-band NIRCam data alone is susceptible to selection of Balmer-break and emission-line interlopers from mid-redshift, and this failure mode appears to become more pernicious at , because of the rarity of these galaxies and because of the more limited number of applicable filters. Using medium-bands F162M, F182M, F210M, F250M, F300M, and F335M will exclude mid-redshift emission lines and allow for detection of more distinctive breaks. We expect that these data will provide robust observational constraints on the number density of galaxies and their rest-UV luminosities, colors, and sizes.
@@
Of course, these exceptionally deep images, up to 46 hours per filter and covering about 9 square arcminutes, will be of high value for the study of intermediate-redshift galaxies as well. To support this, the NIRCam instrument team will conduct NIRCam slitless spectroscopy in Cycle 3 in both the F356W and F444W filters, reaching up to 9 hr of on-source integration time. These will provide spectroscopic redshifts for hundreds of galaxies at with the strong H and/or [OIII] 5007Å lines.
@@
A further opportunity of the Cycle 1 and 2 data is the observation of a very deep NIRSpec MOS pointing on the UDF. The Cycle 1 data was already released in Bunker et al. (2023a); here we describe the science drivers and layout of the Cycle 2 pointing, which will target faint high-redshift galaxies with exposure times up to 104 hrs on-source.
@@
In addition to an overview of the upcoming programs, this paper presents the initial release of NIRCam imaging data and catalogs on this field, already some of the deepest imaging obtained by JWST. The release includes medium-depth flanking fields that connect this footprint to the HUDF region released in the JADES first data release (Rieke et al., 2023).
@@ 2 Cycle 1: Beginning this field as a Coordinated Parallel
The imaging in this field began as the NIRCam Coordinated Parallel of program 1210 (PI: Lützgendorf), which executed the JADES Deep/HST NIRSpec micro-shutter array (MSA) observations in and around the HUDF (Bunker et al., 2023a). Fortuitously, the position angle of these observations put these NIRCam parallels on top of further data from program 1180 (PI: Eisenstein) that included medium-depth NIRCam imaging and deep MIRI F770W imaging, the latter resulting from coordinated parallels of deep NIRCam imaging around the HUDF. These observations are described in Eisenstein et al. (2023). In brief, the 1210 parallel provides 55 open-shutter dual-filter hours of NIRCam imaging in 9 filters: F090W, F115W, F150W, F200W, F277W, F335M, F356W, F410M, and F444W. The 1180 medium imaging overlaps and extends this geometry at flanking depth in all filters save F335M. The MIRI imaging will reach 43 hr depth in F770W in each of 4 adjoining fields.
Further, an assigned observing window in Year 2 has allowed us to place one Coordinated Parallel NIRCam field from the JADES Medium/JWST tier (program 1286, PI: Lützgendorf) in a partially overlapping location about 1 south of the 1210 pointing. This is 11 hours open-shutter of additional data in all the previous nine filters plus F070W. It will add further depth to the JOF imaging, but also will make a deeper region (being the superposition of two medium surveys) that is more square and hence better suited for the NIRSpec MOS footprint. The extra NIRCam imaging also overlaps more of the deep MIRI data.
A simplified view of the GOODS-S region is provided in Figure 1, showing the location of the JOF relative to the HUDF, MUSE, and other JWST programs.
@@ 3.1 Science Motivation: Key Opportunities at
When, how, and why the first stars and black holes formed after the Big Bang is one of the provocative mysteries of the Universe. The galaxies hosting those stars and black holes have remained out of reach until now. With the advent of JWST, we now have a telescope in space that can directly probe and characterize the formation of those first galaxies. This is a key goal of JWST and observational cosmology because the first galaxies probe unknown physics related to star formation (e.g., the efficiency of how gas is converted into stars in early dark matter halos e.g., Behroozi & Silk 2015; Gnedin 2016; Vogelsberger et al. 2020; Kannan et al. 2023; Wilkins et al. 2023) and early structure formation (e.g., Dayal & Ferrara, 2018; Khimey et al., 2021; Lovell et al., 2023; Boylan-Kolchin, 2023). While the low metallicity in this first generation of galaxies should slow H formation and decrease star-formation efficiency (e.g., Krumholz & Dekel, 2012), the gas-rich, compact nature of early galaxies may result in increased gas surface densities and star-formation efficiencies (e.g., Krumholz et al., 2009; Ostriker & Shetty, 2011; Somerville et al., 2015).
@@
There are a wide variety of theoretical predictions for the redshift evolution of the UV luminosity function of galaxies (e.g., Behroozi & Silk, 2015; Mason et al., 2015; Tacchella et al., 2018; Kannan et al., 2023; Wilkins et al., 2023; Katz et al., 2021) These different models typically lead to similar outcomes at regarding number counts and the cosmic star-formation rate density, but they diverge more and more toward higher redshifts due to different treatments of cooling, star formation efficiency, and feedback. Similarly, predictions for the formation and evolution of supermassive black holes vary widely (Habouzit et al., 2022). With JWST, we have the opportunity to push to the redshift frontier of where the physics of these extreme settings cause large differences, factors of 10 and more.
@@
Of course, the redshift frontier offers the opportunity for surprises and more exotic discoveries. As dark matter halos are rapidly growing in this early epoch, galaxies at are a sensitive probe of the interplay of gravitational collapse, gas cooling, and energetic feedback. Indeed, this epoch could even reveal surprises in structure formation and the nature of dark matter: the suppression of small-scale structure in some dark matter models leads to a delayed and subsequently more rapid stellar assembly relative to cold dark matter models (e.g., Gandolfi+ 2022). The exploration of is the most direct way to constrain the onset of cosmic dawn. Together with the constraints on the star-formation efficiency and feedback processes at , JWST can bridge the gap to ongoing 21cm experiments—such as EDGES (Bowman et al., 2018), SARAS 3 (Singh et al., 2022) and REACH (de Lera Acedo et al., 2022)—and thereby creating new synergies to constrain the formation of the first stars and galaxies.
@@
In addition to delivering number density constraints at , JWST observations will constrain the properties of these first galaxies. Specifically, these data are sensitive to the total rest-frame UV luminosity, UV continuum slope , and the morphology of those first structures. The UV continuum slope is a well-known diagnostic for the stellar populations in galaxies (Bouwens et al., 2009; Wilkins et al., 2011; Finkelstein et al., 2012; Bhatawdekar & Conselice, 2021), primarily affected by the dust attenuation and metallicity. Together with the UV luminosity, one can derive dust-corrected star-formation rates (SFRs) and infer the cosmic SFR density at . Furthermore, an estimate of will probe the chemical enrichment and dust physics in the earliest galaxies, extending studies at from lower redshifts (e.g., Ferrara et al., 2022). Very blue UV slopes are a potential signposts of Population III stars or accreting black holes (Schaerer, 2003; Dayal & Ferrara, 2012).
@@
The structure and morphology of galaxies is a well-known tracer of galaxy assembly (e.g., Conselice, 2014). JWST images reveal marginally-resolved galaxies at with compact ( pc) sizes (Robertson et al., 2023; Ono et al., 2023). Deep imaging at will deliver rest-UV size constraints down to a resolution of 300 pc. Measurements of the size and shape (even upper limits) are a sensitive probe of stellar and black hole feedback in those galaxies (e.g., Pillepich et al., 2019; Wu et al., 2020) and could potentially probe the filamentary structure of the gas accretion (e.g., Pandya et al., 2019; Ceverino et al., 2023). Further, one will be sensitive to the multiplicity and close-pair fraction of galaxies at , probing the importance of mergers and star-formation in filaments in these early galaxies (e.g., Wang & Loeb, 2018; O’Leary et al., 2021).

### v2

@@ Abstract
We summarize the properties and initial data release of the JADES Origins Field (JOF), the longest single pointing yet imaged with the James Webb Space Telescope (JWST). This field falls within the GOODS-S region about 8’ south-west of the Hubble Ultra Deep Field (HUDF), where it was formed initially in Cycle 1 as a parallel field of HUDF spectroscopic observations within the JWST Advanced Deep Extragalactic Survey (JADES). This imaging was greatly extended in Cycle 2 program 3215, which observed the JOF for 5 days in six medium-band filters, seeking robust candidates for galaxies. This program also includes ultra-deep parallel NIRSpec spectroscopy (up to 91 hours on-source, summing over the dispersion modes) on the HUDF. Cycle 3 observations from program 4540 added 20 hours of NIRCam slitless spectroscopy and F070W imaging to the JOF. With these three campaigns, the JOF was observed for 380 open-shutter hours with NIRCam using 15 imaging filters and 2 grism bandpasses. Further, parts of the JOF have deep 43 hr MIRI observations in F770W. Taken together, the JOF is one of the most compelling deep fields available with JWST and a powerful window into the early Universe. This paper presents the second data release from JADES, featuring the imaging and catalogs from the year 1 JOF observations.
ma
@@ 1 Introduction
A primary goal of the James Webb Space Telescope (JWST) is to study galaxies to the earliest observable epochs, when the first massive assemblies of gas and stars set in motion the long play of galaxy evolution. Precisely timing when the first galaxies formed after the Big Bang is a key goal of observational cosmology and one of the last great unknowns in our history of the Universe: When and how did the first galaxies form? The primordial star formation and the formation of galaxies in the first 300 Myr (redshift ) after the Big Bang is a sensitive probe of the physics governing the conversion of gas into stars at early times and different dark matter models (e.g., P. S. Behroozi & J. Silk, 2015; P. Dayal & A. Ferrara, 2018; M. Boylan-Kolchin, 2023; R. Kannan et al., 2023). Additionally, these first, unexplored galaxies might harbor surprises such as metal-free stellar populations (Population III) or rapidly accreting young black hole seeds, thereby shedding light onto possible pathways to the formation of the billion-solar-mass black holes observed at (e.g., V. Bromm & N. Yoshida, 2011; H. Xu et al., 2013; A. Smith & V. Bromm, 2019; A. Vikaeus et al., 2022; M. Volonteri et al., 2021; A. Trinca et al., 2023; R. Schneider et al., 2023). Deep fields, both in imaging and associated spectroscopy, are a key opportunity for this work, allowing the astronomy community to coordinate its multi-wavelength resources to best advantage.
@@
Cycle 1 observations with JWST succeeded quickly in pushing the redshift frontier beyond what had been seen with HST. Relying on the complete scattering of photons shortward of Ly 1216Å by the intergalactic medium (IGM), many F115W dropouts at and even some F150W dropouts at were detected as photometric candidates in numerous fields (e.g., M. Castellano et al., 2022; S. L. Finkelstein et al., 2022, 2024; R. P. Naidu et al., 2022a; N. J. Adams et al., 2023; D. Austin et al., 2023; R. Bouwens et al., 2023; C. T. Donnan et al., 2023; Y. Harikane et al., 2023a; P. G. Pérez-González et al., 2023a; B. E. Robertson et al., 2023; L. Whitler et al., 2023; H. Yan et al., 2023; H. Atek et al., 2024; M. Franco et al., 2024; K. N. Hainline et al., 2024a). In the first year, numerous galaxies at , including two at , were spectroscopically confirmed with long NIRSpec exposures (E. Curtis-Lake et al., 2023; K. E. Heintz et al., 2023; T. Morishita & M. Stiavelli, 2023; G. Roberts-Borsani et al., 2023; P. Arrabal Haro et al., 2023a; A. J. Bunker et al., 2023; S. Fujimoto et al., 2023a; R. L. Sanders et al., 2023; M. Tang et al., 2023; B. Wang et al., 2023; C. M. Casey et al., 2024; T. Y.-Y. Hsiao et al., 2024; D. J. McLeod et al., 2024), along with a larger number of galaxies out to where the much brighter [OIII] and H lines are available in the NIRSpec bands.
@@
These early results demonstrated not only JWST’s unrivalled capabilities in successfully exploring the early Universe, but also its potential for pushing the redshift frontier even further. Indeed, it is evident that JWST—with its excellent design and on-orbit performance (J. Rigby et al., 2023)—is fully capable of probing back to even earlier cosmic times. Within this context, we note that the Cycle 1 programs were designed to go deep in F115W and F150W, not in F200W. The few hints of any F200W dropouts are uncertain, potentially mimicked by –5 galaxies with dust-reddened continua and well-placed strong rest-optical emission lines (S. Fujimoto et al., 2023b; R. P. Naidu et al., 2022b; J. A. Zavala et al., 2023; P. Arrabal Haro et al., 2023b; P. G. Pérez-González et al., 2023b).
@@
We describe here the construction of a new deep field, the JADES Origins Field (JOF), designed to push our exploration of early galaxies further than ever before possible. This field was initiated by the JWST Advanced Deep Extragalactic Survey (JADES; D. J. Eisenstein et al., 2023) as JWST NIRCam coordinated parallel observations to a NIRSpec multi-object spectroscopy (MOS) pointing focused on the Hubble Ultra Deep Field (HUDF; S. V. W. Beckwith et al., 2006) within the GOODS-S field (M. Giavalisco et al., 2004). These Cycle 1 observations of JADES were described in D. J. Eisenstein et al. (2023) and A. J. Bunker et al. (2024). New data in Cycle 2 and 3 brought this field to nearly 10 mission days of NIRCam imaging and slitless spectroscopy in addition to over 7 mission days of deep MIRI imaging, making it the deepest JWST imaging yet achieved. The geometry of this field is well suited for future JWST work, as the parallels can span between the JADES Deep mosaic around the HUDF and the JOF at observing windows that have low zodiacal background.
@@
In this paper, we detail the scientific design and implementation of the Cycles 2 and 3 observations. Relative to other deep fields with JWST, this combined data set features very deep medium-band imaging in a total of 15 NIRCam filters. The medium-band imaging is motivated by the goal of having finer sampling of the spectral energy distribution for searches for Ly dropouts at . As we argue in § 3 (and as mentioned above), broad-band NIRCam data alone is susceptible to selection of Balmer-break and emission-line interlopers from mid-redshift, and this failure mode appears to become more pernicious at , because of the rarity of these galaxies and because of the more limited number of applicable filters. Using medium-bands F162M, F182M, F210M, F250M, F300M, and F335M can exclude mid-redshift emission lines and allow for detection of more distinctive breaks. These data provide robust observational constraints on the number density of galaxies and their rest-UV luminosities, colors, and sizes (B. Robertson et al., 2024)
Of course, these exceptionally deep images, up to 46 hours per filter and covering about 9 square arcminutes, are of high value for the study of intermediate-redshift galaxies as well. JADES has conducted 5 NIRSpec MOS pointings on the JOF itself; these data will be presented in JADES Data Release 4 (Curtis-Lake et al., in prep; Scholtz et al., in prep.). An additional pointing using a novel ultra-high multiplex layout for the G235M and G395M grating was taken as part of program 3215 and will be presented in D’Eugenio et al. (in prep.).
@@
To augment this slit spectroscopy, the NIRCam instrument team conducted NIRCam slitless spectroscopy in Cycle 3 in the F322W2, F356W, and F444W filters, reaching up to 9 hr of on-source integration time. These will provide spectroscopic redshifts for hundreds of galaxies at with the strong H and/or [OIII] 5007Å lines.
@@
A further opportunity of the Cycle 1 and 2 data is the observation of a very deep NIRSpec MOS pointing on the UDF. The Cycle 1 data was already released in A. J. Bunker et al. (2024); here we give an overview of the Cycle 2 pointing, which targeted faint high-redshift galaxies with exposure times up to 91 hrs on-source. Further description of the NIRSpec target selection and data properties is presented in JADES Data Release 3 (F. D’Eugenio et al., 2025a).
@@
In addition to an overview of the upcoming programs, this paper presents the initial release of NIRCam imaging data and catalogs on this field, already some of the deepest imaging obtained by JWST. The release includes medium-depth flanking fields that connect this footprint to the HUDF region released in the JADES first data release (M. J. Rieke et al., 2023). Reductions of the deep imaging results will be in an upcoming data release.
@@ 2 Cycle 1: Beginning this field as a Coordinated Parallel
The imaging in this field began as the NIRCam Coordinated Parallel of program 1210 (PI: Lützgendorf), which executed the JADES Deep/HST NIRSpec micro-shutter array (MSA) observations in and around the HUDF (A. J. Bunker et al., 2024). Fortuitously, the position angle of these observations put these NIRCam parallels on top of further data from program 1180 (PI: Eisenstein) that included medium-depth NIRCam imaging and deep MIRI F770W imaging, the latter resulting from coordinated parallels of deep NIRCam imaging around the HUDF. These observations are described in D. J. Eisenstein et al. (2023). In brief, the 1210 parallel provides 55 open-shutter dual-filter hours of NIRCam imaging in 9 filters: F090W, F115W, F150W, F200W, F277W, F335M, F356W, F410M, and F444W. The 1180 medium imaging overlaps and extends this geometry at flanking depth in all filters save F335M. The MIRI imaging reach 43 hr depth in F770W in each of 4 adjoining fields.
Further, an assigned observing window in Year 2 allowed us to place one Coordinated Parallel NIRCam field from the JADES Medium/JWST tier (program 1286, PI: Lützgendorf) in a partially overlapping location about 1′ south of the 1210 pointing. This is 11 hours open-shutter of additional data in all the previous nine filters plus F070W. It adds further depth to the JOF imaging, but also makes a deeper region (being the superposition of two medium surveys) that is more square and hence better suited for the NIRSpec MOS footprint. The extra NIRCam imaging also overlaps more of the deep MIRI data.
A simplified view of the GOODS-S region is provided in Figure 1, showing the location of the JOF relative to the HUDF, CANDELS, and other JWST programs.
@@ 3.1 Science Motivation: Key Opportunities at
When, how, and why the first stars and black holes formed after the Big Bang is one of the provocative mysteries of the Universe. The galaxies hosting those stars and black holes have remained out of reach until now. With the advent of JWST, we now have a telescope in space that can directly probe and characterize the formation of those first galaxies. This is a key goal of JWST and observational cosmology because the first galaxies probe unknown physics related to star formation (e.g., the efficiency of how gas is converted into stars in early dark matter halos e.g., P. S. Behroozi & J. Silk 2015; N. Y. Gnedin 2016; M. Vogelsberger et al. 2020; R. Kannan et al. 2023; S. M. Wilkins et al. 2023) and early structure formation (e.g., P. Dayal & A. Ferrara, 2018; D. Khimey et al., 2021; C. C. Lovell et al., 2023; M. Boylan-Kolchin, 2023). While the low metallicity in this first generation of galaxies should slow H2 formation and decrease star-formation efficiency (e.g., M. R. Krumholz & A. Dekel, 2012), the gas-rich, compact nature of early galaxies may result in increased gas surface densities and star-formation efficiencies (e.g., M. R. Krumholz et al., 2009; E. C. Ostriker & R. Shetty, 2011; R. S. Somerville et al., 2015).
@@
There are a wide variety of pre-flight theoretical predictions for the redshift evolution of the UV luminosity function of galaxies (e.g., P. S. Behroozi & J. Silk, 2015; C. A. Mason et al., 2015; S. Tacchella et al., 2018; R. Kannan et al., 2023; S. M. Wilkins et al., 2023; H. Katz et al., 2021) These different models typically lead to similar outcomes at regarding number counts and the cosmic star-formation rate density, but they diverge more and more toward higher redshifts due to different treatments of cooling, star formation efficiency, and feedback. Similarly, predictions for the formation and evolution of supermassive black holes vary widely (M. Habouzit et al., 2022). With JWST, we have the opportunity to push to the redshift frontier of where the physics of these extreme settings cause large differences, factors of 10 and more.
@@
Of course, the redshift frontier offers the opportunity for surprises and more exotic discoveries. As dark matter halos are rapidly growing in this early epoch, galaxies at are a sensitive probe of the interplay of gravitational collapse, gas cooling, and energetic feedback. Indeed, this epoch could even reveal surprises in structure formation and the nature of dark matter: the suppression of small-scale structure in some dark matter models leads to a delayed and subsequently more rapid stellar assembly relative to cold dark matter models (e.g., Gandolfi+ 2022). The exploration of is the most direct way to constrain the onset of cosmic dawn. Together with the constraints on the star-formation efficiency and feedback processes at , JWST can bridge the gap to ongoing 21cm experiments—such as EDGES (J. D. Bowman et al., 2018), SARAS 3 (S. Singh et al., 2022) and REACH (E. de Lera Acedo et al., 2022)—and thereby creating new synergies to constrain the formation of the first stars and galaxies.
@@
In addition to delivering number density constraints at , JWST observations constrain the properties of these first galaxies. Specifically, these data are sensitive to the total rest-frame UV luminosity, UV continuum slope , and the morphology of those first structures. The UV continuum slope is a well-known diagnostic for the stellar populations in galaxies (R. J. Bouwens et al., 2009; S. M. Wilkins et al., 2011; S. L. Finkelstein et al., 2012; R. Bhatawdekar & C. J. Conselice, 2021), primarily affected by the dust attenuation and metallicity. Together with the UV luminosity, one can derive dust-corrected star-formation rates (SFRs) and infer the cosmic SFR density at . Furthermore, an estimate of probes the chemical enrichment and dust physics in the earliest galaxies, extending studies at from lower redshifts (e.g., A. Ferrara et al., 2022). Very blue UV slopes are a potential signposts of Population III stars or accreting black holes (D. Schaerer, 2003; P. Dayal & A. Ferrara, 2012).
@@
The structure and morphology of galaxies is a well-known tracer of galaxy assembly (e.g., C. J. Conselice, 2014). JWST images reveal marginally-resolved galaxies at with compact ( pc) sizes (B. E. Robertson et al., 2023; Y. Ono et al., 2023). Deep imaging at can deliver rest-UV size constraints down to a resolution of 300 pc. Measurements of the size and shape (even upper limits) are a sensitive probe of stellar and black hole feedback in those galaxies (e.g., A. Pillepich et al., 2019; X. Wu et al., 2020) and could potentially probe the filamentary structure of the gas accretion (e.g., V. Pandya et al., 2019; D. Ceverino et al., 2023). Further, one can be sensitive to the multiplicity and close-pair fraction of galaxies at , probing the importance of mergers and star-formation in filaments in these early galaxies (e.g., X. Wang & A. Loeb, 2018; J. A. O’Leary et al., 2021).